# Spatial vertical regression for spatial panel data: Evaluating the effect of the Florentine tramway’s first line on commercial vitality

### v2

@@ 2 Illustration of simulation specifications
3 Additional information on our study and results
3.1 De-trended outcome time series
@@
3.2 Point estimates across treated units
@@
3.3 MCMC diagnostics
@@
Figure S.8 reports the traceplots and posterior distributions for the parameters estimated in the application. We cannot notice convergence problems using N=15000 iterations. Note that the posterior distribution of $w$, the $[0,1]$ parameter controlling the radio between spatial and IID errors is shifted toward 1, suggesting the presence of spatially correlated errors.
Figure S.8: Traceplot and Posterior distribution for the hyperparaters of the model ($\rho^{2}_{e},\rho^{2}_{b},\sigma^{2}_{e},\sigma^{2}_{b},w$).
3.4 Point estimates from all methods compared in our simulations
In Figure S.9 we compare the point estimates for the different methods we propose in the simulations. We notice that all methods provide similar results, leading to a similar interpretation of results.
@@ 3.4 Point estimates from all methods compared in our simulations
Figure S.9: Comparison across alternative methods between posterior medians of the treatment effects $\Delta_{it}$ for areas within $d-$seconds of walking from a tramway stop, $d=120,240,360,480,600$. First vertical line: tramway worksite starts (2006) - Second vertical line: tramway operational (2010).
4 Additional simulation results
4.1 Figures of bias and MSE across all simulations
@@
In this subsection we provide the graphs for bias and MSE under the several scenarios we tested. In particular Figures S.10, S.11, S.12 reports the average bias, and Figures S.13, S.14, S.15 the MSE for $T_{0}=10$, $T_{0}=20$, and $T_{0}=40$, respectively. In all figures, the rows correspond to the different lengthscale parameter employed $\rho_{s}^{2}=\{0.001^{2},0.2^{2},0.4^{2},0.6^{2}\}$, while the columns represent the configurations for the error terms.
@@
For the shortest pre-intervention time period ($T_{0}=10$), the various estimation methods considered are still able to provide reasonably accurate predictions of the outcome (Figure S.10). The only model that consistently underestimates the counterfactual is OLS, which is known to be prone to overfitting when the pre-treatment period is short. In terms of MSE (Figure S.13), OLS yields the highest post-treatment MSE, reflecting the same issues of overfitting. SVR achieves a lower MSE than its counterpart without spatially correlated priors (BVR), but a higher MSE compared to methods with stronger penalization or numerical constraints on the coefficients (SR, BSC, SC). Notably, even in this configuration, SVR is able to leverage spatial information through the priors and reduce the MSE as the lengthscale increases.
@@
For longer pre-treatment time periods increases ($T_{0}\in\{20,40\}$), the accuracy of the counterfactual estimates remains high across all methods (see Figures S.11 and S.12). In scenarios characterized by low variance and uncorrelated errors, methods based on unconstrained models (BVR, SVR, OLS, SR) yield very similar and more stable results compared to those relying on constrained coefficients (BSC, SC). In terms of MSE (see Figures S.14 and S.15), OLS remains the noisiest estimator in terms of coefficient variability for $T_{0}=20$, although its performance becomes more comparable to that of other methods for $T_{0}=40$. SVR consistently delivers the lowest post-treatment MSE by effectively exploiting the spatial structure encoded in the priors. We notice that the MSE decreases with increasing lengthscale–see, for instance, column 3 in Figure S.15.
@@
Figures S.17, S.17, and S.18 show the MSE of methods that do not explicitly leverage spatial information (SC, BVR, and BSC) as the spatial structure in the control units’ coefficients increases, across different lengths of the pre-treatment period and model error configurations. The corresponding figure for SVR is given in Figure 3. It is evident that SVR reduces the MSE as spatial correlation increases, whereas all other methods show no substantial change in performance under the same conditions, except perhaps for BSC in the short pre-treatment period ($T_{0}=10$). This highlights the source of the performance improvement in the model we propose—namely, its ability to incorporate and exploit spatial structure effectively.
@@
4.2 Additional tables of simulation results
The tables presented provide a detailed analysis of the performance of several statistical methodologies for estimating counterfactual outcomes, focusing on bias, Mean Squared Error (MSE), and coverage intervals. Tables S.1, S.2 in the Supplement, Table 1 in Section 6, and Table S.3 report the bias of all approaches under different strengths of spatial structure, $\rho^{2}_{s}\in\{0.001^{2},0.2^{2},0.4^{2},0.6^{2}\}$, respectively. Similarly, Tables S.4 and S.5 in the Supplement, Table 2 in Section 6, and Table S.6 in the Supplement include the MSE, and Tables S.7, S.8 in the Supplement, Table 3 in Section 6, and S.9 in the Supplement include the coverage across values of $\rho^{2}_{s}$. These tables repeat some of the information discussed using the figures in Supplement 4.1.
OLS consistently shows strong negative bias for short pre-treatment periods ($T_{0}=10$), due to overfitting. SVR stands out for its ability to reduce bias as spatial correlation increases, particularly for longer pre-treatment periods. In terms of MSE, OLS performs poorly for low $T_{0}$, with very high error values. SVR, on the other hand, shows substantial improvements as spatial correlation increases, particularly when $T_{0}=40$. In contrast, SC, BSC, and BVR show more consistent MSE regardless of spatial dependence, indicating less responsiveness to spatial information. Coverage analysis further highlights OLS’s limitations, with near-zero interval coverage for small $T_{0}$. BVR and SVR, however, achieve close-to-nominal 95% coverage, even under spatial noise. SVR maintains reliable uncertainty estimates across varying conditions, underlining its robustness. Overall, SVR consistently reduces bias and MSE while maintaining accurate uncertainty quantification, particularly in spatially dependent settings. This underscores the value of incorporating spatial structure into counterfactual estimation models.
@@ 4.2 Additional tables of simulation results
SVR 0.00 0.00 -0.01
Table S.1: Average post-treatment bias of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Matrix Completion (SVR) in scenarios with spatially non-correlated weights ($\rho^{2}_{s}=0.001^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR -0.01 -0.01 -0.01
Table S.2: Average post-treatment bias of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Matrix Completion (SVR) in scenarios with spatially low-correlated weights ($\rho^{2}_{s}=0.2^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.00 0.00 0.00
Table S.3: Average post-treatment bias of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Matrix Completion (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.6^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.44 0.47 0.79
Table S.4: Average post-treatment MSE of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Matrix Completion (SVR) in scenarios with spatially non-correlated weights ($\rho^{2}_{s}=0.001^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.37 0.45 0.75
Table S.5: Average post-treatment MSE of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Matrix Completion (SVR) in scenarios with spatially low-correlated weights ($\rho^{2}_{s}=0.2^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.32 0.39 0.66
Table S.6: Average post-treatment MSE of Synthetic Control (SC); Separable Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Matrix Completion (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.6^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.95 0.94 0.94
Table S.7: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Bayesian Spatial Matrix Completion (SVR) in scenarios with spatially non-correlated weights ($\rho^{2}_{s}=0.001^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Table S.8: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Spatial Vertical Regression (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.2^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Table S.9: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BVR) and Spatial Vertical Regression (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.6^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.

### v3

@@ Appendix B Illustration of simulation specifications
Appendix C Additional information on our study and results
C.1 De-trended outcome time series
@@
C.2 Point estimates across treated units
@@
C.3 Additional results for stores selling durable goods and non durable goods
The impact of a tramway construction may have different effects if we consider different types of stores. Grossi et al. [2025] distinguish two different types of stores, according to the nature of goods provided:
  • Stores that sell durable goods, such as electronics, clothes, furniture. The purchase frequency for these types of goods is low, hereafter we can consider that the customers’ choices may be less affected by the new tramway line, both in the construction phase, and in the operational period of the tramway.
  • Stores that are selling non durable goods: such as restaurant, bars, news kiosk, book shops, grocery stores, bakeries, etc . These stores have a high-frequency of purchase and thus, may be more affected by the construction of the tramway.
Figures S.10, S.11 and table Table S.2 reports the estimation results at different distances from the closest tramway stops for stores that sells durable goods. For these stores we can see a positive effect for the areas located at higher distance from the tramway stops, with credibility intervals that do not comprise zero for the last time periods (2013-2016).
Figures S.12, S.13 and table Table S.3 reports the same results for stores that are selling non-durable goods. In this case we can underline a positive effect after the construction of the tramway, after a first contraction during the construction period, which is more evident for the areas between 120 and 480 seconds walking from the tramway.
C.4 MCMC diagnostics
@@
Figure S.8 presents the traceplots and the histograms of the posterior distributions for the parameters $\rho^{2}_{e}$, $\rho^{2}_{b}$, $\sigma^{2}_{e}$ and $w$ estimated using 10,000 iterations (5,000 warmup) across three chains. Figure S.8 illustrates that the traceplots for all monitored parameters exhibit adequate mixing, with no visible trends or sticking behavior across chains. Consistently $\widehat{R}$ values are close to 1 and Stan reports no warning for the model fit. These diagnostics provide evidence that the sampler has achieved satisfactory convergence and that posterior inference is reliable. Note that the posterior distribution of $w$, the $[0,1]$ parameter controlling the radio between spatial and IID errors is shifted toward 1, suggesting the presence of spatially correlated errors.
Figure S.8: Traceplot and Posterior distribution for the hyperparaters of the model ($\rho^{2}_{e},\rho^{2}_{b},\sigma^{2}_{e},w$).
C.5 Point estimates from all methods compared in our simulations
In Figure S.9 and Table S.1 we compare the point estimates for the different methods we propose in the simulations. We notice that all methods provide similar results, leading to a similar interpretation of results. Table S.1 reports RMSE, and average treatment effect for the served areas during the construction of the tramway and the operational period of the tramway. Table S.1 shows that differences across methods mainly concern pre-treatment fit and uncertainty quantification. Across most approaches, estimates point to moderate positive effects during the construction phase and more heterogeneous effects during the operational phase. Methods achieving very low pre-treatment RMSPE, such as OLS and BVR, tend to suffer from overfitting and inferential problems, as evidenced by degenerate or excessively wide confidence intervals. In contrast, SVR exhibits an intermediate pre-treatment RMSPE, higher than BVR but substantially lower than SC, BSC, and SR, while providing coherent and stable uncertainty quantification.
@@ C.5 Point estimates from all methods compared in our simulations
Figure S.9: Comparison across alternative methods between posterior medians of the treatment effects $\Delta_{it}$ for areas within $d$-seconds of walking from a tramway stop, $d=120,240,360,480,600$. The two vertical lines represent the beginning of the tramways construction and operational phases in 2006 and 2010, respectively.
Table S.1: Estimates, credible intervals, and pre-treatment RMSPE for all methods.
Method Period Type 120 240 360 480 600
SVR Construction Estimate 3.306 1.064 3.453 6.232 6.356
SVR Construction CI (-1.003, 7.572) (-5.596, 7.670) (-6.772, 14.028) (-3.239, 16.185) (-9.463, 22.689)
SVR Operational Estimate 1.803 0.359 -2.510 3.228 3.650
SVR Operational CI (-2.938, 6.485) (-7.076, 7.839) (-14.645, 10.008) (-8.065, 15.378) (-14.930, 23.361)
SVR Pre-treatment RMSPE 0.834 1.584 1.291 2.094 2.769
BVR Construction Estimate 2.547 1.551 -1.253 2.933 0.341
BVR Construction CI (-3.315, 8.491) (-9.669, 12.214) (-13.112, 12.205) (-9.927, 17.441) (-21.151, 25.106)
BVR Operational Estimate -0.181 -4.222 -2.573 -4.722 1.401
BVR Operational CI (-5.650, 7.504) (-15.575, 10.434) (-15.850, 11.561) (-17.910, 13.760) (-23.427, 26.346)
BVR Pre-treatment RMSPE 0.066 0.569 0.085 0.312 0.288
SC Construction Estimate 2.439 2.799 3.643 7.608 8.916
SC Construction CI
SC Operational Estimate 1.744 1.124 0.147 6.606 6.058
SC Operational CI
SC Pre-treatment RMSPE 1.314 1.673 1.767 3.178 3.798
BSC Construction Estimate 3.435 0.803 3.843 6.851 6.329
BSC Construction CI (-1.094, 7.963) (-4.751, 6.260) (-2.617, 10.465) (-2.477, 16.192) (-6.479, 18.821)
BSC Operational Estimate 0.669 1.186 2.126 11.291 15.525
BSC Operational CI (-3.976, 5.170) (-4.559, 6.809) (-4.848, 9.159) (1.784, 21.111) (2.575, 29.385)
BSC Pre-treatment RMSPE 2.069 2.320 2.505 4.166 5.254
SR Construction Estimate 3.654 1.012 3.778 7.173 6.833
SR Construction CI
SR Operational Estimate 0.923 0.940 1.527 9.272 13.747
SR Operational CI
SR Pre-treatment RMSPE 1.732 2.174 2.292 3.630 4.837
OLS Construction Estimate 2.574 2.643 -2.034 3.131 -1.021
OLS Construction CI (2.574, 2.574) (2.643, 2.643) (-2.034, -2.034) (3.131, 3.131) (-1.021, -1.021)
OLS Operational Estimate -0.910 -9.288 -2.758 -7.560 1.696
OLS Operational CI (-0.910, -0.910) (-9.288, -9.288) (-2.758, -2.758) (-7.560, -7.560) (1.696, 1.696)
OLS Pre-treatment RMSPE 0.000 0.000 0.000 0.000 0.000
Figure S.10: Posterior medians (lines) and 95% posterior credible intervals (shaded areas) of the treatment effects $\Delta_{it}$ for areas within $d$-seconds of walking from a tramway stop, $d=120,240,360,480,600$. The two vertical lines represent the beginning of the tramways construction and operational phases in 2006 and 2010, respectively.
Figure S.11: Comparison across alternative methods between posterior medians of the treatment effects $\Delta_{it}$ for areas within $d$-seconds of walking from a tramway stop, $d=120,240,360,480,600$. The two vertical lines represent the beginning of the tramways construction and operational phases in 2006 and 2010, respectively.
Figure S.12: Posterior medians (lines) and 95% posterior credible intervals (shaded areas) of the treatment effects $\Delta_{it}$ for areas within $d$-seconds of walking from a tramway stop, $d=120,240,360,480,600$. The two vertical lines represent the beginning of the tramways construction and operational phases in 2006 and 2010, respectively.
Figure S.13: Comparison across alternative methods between posterior medians of the treatment effects $\Delta_{it}$ for areas within $d$-seconds of walking from a tramway stop, $d=120,240,360,480,600$. The two vertical lines represent the beginning of the tramways construction and operational phases in 2006 and 2010, respectively.
Table S.2: Posterior medians and 95% posterior credible intervals of the intertemporal average causal effects $\Delta_{i}$ for stores selling durable goods, for areas within $d$-seconds walking distance from the tramway stop, $d=120,240,360,480,600$, and posterior median of the pre-treatment root mean squared prediction error (RMSPE).
120 240 360 480 600
Construction period
Estimate 1.197 4.028 5.963 9.072 8.345
CI (-3.239, 5.485) (-3.348, 11.379) (-4.245, 16.294) (-1.070, 19.452) (-2.431, 19.546)
Operational period
Estimate -0.956 3.128 6.703 11.390 13.212
CI (-6.697, 4.627) (-6.362, 12.553) (-6.394, 19.948) (-1.459, 24.805) (-0.552, 27.751)
RMSPE 1.103 1.643 1.570 1.817 1.955
Table S.3: Posterior medians and 95% posterior credible intervals of the intertemporal average causal effects $\Delta_{i}$ for stores selling non durable goods, for areas within $d$-seconds walking distance from the tramway stop, $d=120,240,360,480,600$, and posterior median of the pre-treatment root mean squared prediction error (RMSPE).
120 240 360 480 600
Construction period
Estimate 2.012 -4.101 -3.341 -2.979 -1.729
CI (0.222, 3.743) (-6.841, -1.371) (-9.281, 2.670) (-7.884, 2.042) (-14.420, 11.462)
Operational period
Estimate 0.883 -1.493 -2.974 3.756 4.695
CI (-1.434, 3.137) (-5.018, 2.008) (-10.595, 4.732) (-2.459, 10.245) (-11.515, 21.817)
RMSPE 1.262 1.731 2.551 2.453 3.326
Appendix D Additional simulation results
D.1 Figures of bias and MSE across all simulations
@@
In this subsection we provide the graphs for bias and MSE under the several scenarios we tested. In particular Figures S.14, S.15, S.16 reports the average bias, and Figures S.17, S.18, S.19 the MSE for $T_{0}=10$, $T_{0}=20$, and $T_{0}=40$, respectively. In all figures, the rows correspond to the different lengthscale parameter employed $\rho_{s}^{2}=\{0.001^{2},0.2^{2},0.4^{2},0.6^{2}\}$, while the columns represent the configurations for the error terms.
@@
For the shortest pre-intervention time period ($T_{0}=10$), all estimation methods considered, except OLS, are still able to provide reasonably accurate predictions of the outcome (Figure S.14). In contrast, OLS, which is known to be prone to overfitting when the pre-treatment period is short, consistently underestimates the counterfactual. In terms of MSE (Figure S.17), OLS yields the highest post-treatment MSE, reflecting the same issues of overfitting. SVR achieves the lowest MSE across all methods when the errors are IID under all the scenarios considered, and also with spatially correlated residuals, when the time series of treated units are correlated, which is a reasonably common situation. SR has slightly lower MSE when the post treatment is very long, and the pre-treatment is short, but the estimates are comparable. Notably, SVR is able to leverage spatial information through the priors and reduce the MSE as the lengthscale increases, confirming the implicit regularization that we exploit with gaussian process priors.
@@
When pre-treatment length increases ($T_{0}\in\{20,40\}$), the accuracy of the counterfactual estimates remains high across all methods (see Figures S.15 and S.16). In scenarios characterized by low variance and uncorrelated errors, methods based on unconstrained models (BVR, SVR, OLS, SR) yield very similar and more stable results compared to those relying on constrained coefficients (BSC, SC). In terms of MSE (see Figures S.18 and S.19), OLS remains the noisiest estimator in terms of coefficient variability for $T_{0}=20$, although its performance becomes more comparable to that of other methods for $T_{0}=40$. SVR consistently delivers the lowest post-treatment MSE by effectively exploiting the spatial structure encoded in the priors. We notice that the MSE decreases with increasing lengthscale–see, for instance, column 3 in Figure S.19.
@@
Figures S.21, S.21, and S.22 show the MSE of methods that do not explicitly leverage spatial information (SC, BVR, and BSC) as the spatial structure in the control units’ coefficients increases, across different lengths of the pre-treatment period and model error configurations. The corresponding figure for SVR is given in Figure 3. It is evident that SVR reduces the MSE as spatial correlation increases, whereas all other methods show no substantial change in performance under the same conditions. This highlights the source of the performance improvement in the model we propose—namely, its ability to incorporate and exploit spatial structure effectively.
@@
D.2 Additional tables of simulation results
The tables presented provide a detailed analysis of the performance of several statistical methodologies for estimating counterfactual outcomes, focusing on bias, Mean Squared Error (MSE), and coverage intervals. Tables S.4, S.5 in the Supplement, Table 1 in Section 6, and Table S.6 report the bias of all approaches under different strengths of spatial structure, $\rho^{2}_{s}\in\{0.001^{2},0.2^{2},0.4^{2},0.6^{2}\}$, respectively. Similarly, Tables S.7 and S.8 in the Supplement, Table 2 in Section 6, and Table S.9 in the Supplement include the MSE, and Tables S.10, S.11 in the Supplement, Table 3 in Section 6, and S.12 in the Supplement include the coverage across values of $\rho^{2}_{s}$. These tables repeat some of the information discussed using the figures in Supplement D.1.
SVR stands out for its ability to reduce bias as spatial correlation increases, particularly for longer pre-treatment periods. In terms of MSE, OLS performs poorly for low $T_{0}$, with very high error values. SVR, on the other hand, shows substantial improvements as spatial correlation increases, particularly when $T_{0}=40$. Coverage analysis further highlights OLS’s limitations, with near-zero interval coverage for small $T_{0}$. BVR and SVR, however, achieve close-to-nominal 95% coverage, even under spatial noise. SVR maintains reliable uncertainty estimates across varying conditions, underlining its robustness. BSC tends to undercover the 95% nominal confidence interval. Overall, SVR consistently reduces bias and MSE while maintaining accurate uncertainty quantification, particularly in spatially dependent settings. This underscores the value of incorporating spatial structure into counterfactual estimation models.
@@ D.2 Additional tables of simulation results
SVR 0.00 0.00 0.00
Table S.4: Average post-treatment bias of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially non-correlated weights ($\rho^{2}_{s}=0.001^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods. Bias and outcome mean value are expressed in the same unit of measure.
@@
SVR -0.01 00.00 -0.01
Table S.5: Average post-treatment bias of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially low-correlated weights ($\rho^{2}_{s}=0.2^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods. Bias and outcome mean value are expressed in the same unit of measure.
@@
Table S.6: Average post-treatment bias of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.6^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods. Bias and outcome mean value are expressed in the same unit of measure.
@@
SVR 0.43 0.46 0.74
Table S.7: Average post-treatment MSE of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially non-correlated weights ($\rho^{2}_{s}=0.001^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.37 0.44 0.72
Table S.8: Average post-treatment MSE of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially low-correlated weights ($\rho^{2}_{s}=0.2^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.32 0.38 0.64
Table S.9: Average post-treatment MSE of Synthetic Control (SC); Separate Ridge (SR) regression; OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.6^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
SVR 0.94 0.94 0.94
Table S.10: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially non-correlated weights ($\rho^{2}_{s}=0.001^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Table S.11: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.2^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.
@@
Table S.12: Average post-treatment coverage of 95% confidence / posterior credible intervals for the imputed post-treatement control potential outcomes for the treated units based on OLS vertical regression; Bayesian Vertical Regression (BVR); Bayesian Synthetic Control (BSC) and Spatial Vertical Regression (SVR) in scenarios with spatially high-correlated weights ($\rho^{2}_{s}=0.6^{2}$). Columns correspond to the length of the pre-intervention period $T_{0}$ and the three cases of residual error structure, IID, Spatial with 40% noise, and Spatial with 70% noise. Rows correspond to the length of the post-intervention period and the different methods.